# Towards General Industrial Intelligence: A Survey on Industrial IoT-Enhanced Continual Large Models

### v1

@@ II-C2 Detection/Localization
Industrial Anomaly Detection. Industrial anomaly detection is geared towards identifying behaviours that deviate from normal operational patterns through continuous production process monitoring [14]. Popular datasets such as MVTec-AD [57] and VisA [58] are extensively used for this purpose. MVTec-AD specifically focuses on surface defect detection across various industrial objects made of different materials and shapes. VisA provides a broader spectrum of visual anomalies, enabling extensive testing of models’ generalization capabilities in complex scenarios. Deploying deep learning techniques on these datasets has demonstrated strong potential in effectively addressing unknown types of anomalies (see Fig. 8).
@@ II-C3 Planning & Control
Autonomous Driving. Autonomous driving services depends on active sensing from modules such as camera, LiDAR, radar, and communication units [31]. Autonomous driving perception and motion planning integrate generative LLMs with path planning to enable natural language control of autonomous vehicles (see Fig. 11). The process begins with speech-to-text technology converting verbal descriptions into text, which LLMs then interpret to generate executable path planning code. This code is tested in simulation environments, providing feedback for optimizing vehicle operation, enhancing adaptability, and decision-making accuracy [65, 66].
@@
Figure 11: LLM planning and control for autonomous driving.
Co-Design of Communication and Control. As a critical aspect of building the IIoT, the design of communication and control systems has always been a major focus[74]. Due to the strong coupling relationship between the two, it is crucial to develop new metrics that can simultaneously evaluate the performance of both communication and control systems [75]. Wang et al. [76] used the Age of Information (AoI) to characterize “information freshness” and studied the impact of average AoI on control costs. By minimizing AoI, they aimed to reduce the combined costs of control and communication. Chen et al. [77] considered scenarios with multiple users, addressing both uplink and downlink wireless transmissions. Their work integrated multiple antennas on the remote end to improve signal reception quality and developed a loop information age metric. This metric not only captures information freshness but also dynamically allocates the rounds of transmission attempts based on the success or failure of UL and DL transmissions, thereby enhancing communication reliability while ensuring control convergence speed.
@@ II-C4 General Q&A
In the industrial sector, QA systems are crucial for addressing complex queries, supporting decision-making, and improving operational efficiency. By integrating into enterprise knowledge management systems, QA technology helps engineers and operators quickly find solutions, reducing downtime and optimizing maintenance processes.
@@
While LLMs perform well in open-domain QA tasks, their application in specific industrial domains presents challenges. Retrieval-based LLMs often struggle with complex industrial queries requiring deep domain-specific knowledge. Improvement strategies include new interaction paradigms that integrate domain documentation and fine-tuning techniques to provide domain-specific knowledge during runtime (see Fig. 13). Additionally, specialized QA datasets for industrial applications, such as the 32k QA pair dataset in the cloud domain by Wang et al. [79], offer new benchmarks for evaluating and improving LLMs.
Figure 12: Communication and control system.
III Large Models for General Industrial Intelligence
LMs, owing to their extensive internal knowledge and reasoning capabilities, enhance GII in the following two fundamental ways:
(1) As effective enhancers of data, LMs leverage their extensive external knowledge and analytical capabilities to improve the quality of industrial data and optimize existing methodologies;
(2) As efficient predictor in industrial tasks, LMs utilize their rich internal knowledge and emerging inferential abilities to effectively handle various predictive tasks, including both tuning and non-tuning predictors based on whether access to LMs parameters.
III-A LMs as a Data Enhancer
III-A1 Data Denoise
@@ III-A2 Data Selection
Figure 13: LLMs for industrial question answering.
III-A3 Data Augmentation
The potential of LLMs in data augmentation is prominent, particularly in interpreting industrial time series data through textual descriptions and summaries. SignalGPT, specifically trained for diagnosing electrocardiogram signals, exemplifies this adaptability [85]. Zeng et al. [86] demonstrate the use of Textual Entailment models in zero-shot event detection, enhancing performance through initial large-scale text annotation and subsequent model fine-tuning.
LLMs also excel in integrating diverse industrial data sources, enhancing data processing robustness and decision-making. For instance, Hou et al. [87] leverage GPT-4 to refine hierarchical feature definitions, reducing manual annotation needs and bolstering model stability during testing (see Fig. 14).
III-A4 Data Generation
@@
Furthermore, we have included a comparison between raw and generated data, as illustrated in Fig. 15, highlighting the ability of DDPM to produce samples closely resembling the original time series data.
Figure 14: The generation of unique domain descriptions.
@@
III-A5 Reprogramming
Discussion. LMs, effectively mitigate the sparsity and noise inherent in industrial data as data enhancers, endowing existing general industrial intelligence models with enriched external knowledge and augmented analytical capabilities. This technology’s plug-and-play nature offers flexibility in addressing real-world challenges associated with industrial data and models. Nonetheless, a significant barrier to utilizing LMs as enhancers is the substantial time and cost of processing large-scale datasets. Moreover, the inherent diversity of industrial data and the breadth of application scenarios add complexity to creating universally effective enhancers for various contexts.
LMs as data enhancers represent a promising new direction for augmenting industrial data and models, meriting further exploration in future research. Subsequent studies should concentrate on developing efficient, dependable, and universally applicable plug-and-play solutions that address real-world issues like data sparsity and noise while considering the time and cost efficiency for large-scale dataset applications.
III-B LMs as a Tuning Predictor

### v2

@@ II-C2 Detection/Localization
Industrial Anomaly Detection. As illustrated in Fig. 6, industrial anomaly detection aims to identify behaviours that deviate from normal operational patterns through continuous monitoring of production processes [53]. Popular datasets such as MVTec-AD [54] and VisA [55] are extensively used to benchmark detection models. MVTec-AD focuses on surface defect detection across various industrial objects with diverse materials and shapes, while VisA offers a broader range of visual anomalies, facilitating comprehensive evaluation of model generalization capabilities in complex scenarios.
@@ II-C3 Planning & Control
Autonomous Driving. Autonomous driving services depends on active sensing from modules such as camera, LiDAR, radar, and communication units [62]. As shown in Fig. 8, integrating LLMs with the IIoT offers great potential for enhancing vehicle personalization and adaptability in autonomous driving, particularly in open-world scenarios [63, 64]. The process begins with speech-to-text technology converting verbal descriptions into text, which LLMs then interpret to generate executable path planning code. This code is tested in simulation environments, providing feedback for optimizing vehicle operation, enhancing adaptability, and decision-making accuracy [65, 66].
Figure 8: LLM planning and control for autonomous driving.
@@
Fig. 8 illustrates the application of large models in autonomous driving scenarios. The first row depicts the situation without a large model, where the ego vehicle, despite decelerating, maintains a relatively high speed when encountering a pedestrian crossing, ultimately resulting in a collision. In contrast, the second row demonstrates the scenario when a “stop” command is issued. The ego vehicle promptly executes a braking response, reducing its speed from 10.65 m/s to 1.06 m/s within just 2 seconds, successfully avoiding the collision.
@@
Co-Design of Communication and Control. As a critical aspect of building the IIoT, the design of communication and control systems has always been a major focus[74]. Due to the strong coupling relationship between the two, it is crucial to develop new metrics that can simultaneously evaluate the performance of both communication and control systems [75]. Wang et al. [76] used the Age of Information (AoI) to characterize “information freshness” and studied the impact of average AoI on control costs. By minimizing AoI, they aimed to reduce the combined costs of control and communication. Chen et al. [77, 78] considered scenarios with multiple users, addressing both uplink and downlink wireless transmissions. Their work integrated multiple antennas on the remote end to improve signal reception quality and developed a loop information age metric. This metric not only captures information freshness but also dynamically allocates the rounds of transmission attempts based on the success or failure of UL and DL transmissions, thereby enhancing communication reliability while ensuring control convergence speed.
Figure 10: Communication and control system [78].
@@ II-C4 General Q&A
In the industrial sector, QA systems play a crucial role in addressing complex queries, supporting decision-making, and improving operational efficiency. As illustrated in Fig. 11, these systems integrate into enterprise knowledge management frameworks to assist engineers and operators in quickly finding solutions, thereby reducing downtime and optimizing maintenance processes.
@@
While LLMs perform well in open-domain QA tasks, their application in specific industrial domains presents unique challenges. Retrieval-based LLMs often struggle with complex industrial queries requiring deep domain-specific knowledge. Strategies to address these limitations include novel interaction paradigms that incorporate domain documentation and fine-tuning techniques to provide relevant knowledge during runtime. Furthermore, specialized QA datasets for industrial applications, such as the 32k QA pair dataset in the cloud domain by Wang et al. [80], establish new benchmarks for evaluating and improving LLM performance in industrial contexts.
II-D LMs Evaluation
To effectively evaluate the performance of LLMs, a common approach is to select different capability dimensions and construct corresponding evaluation tasks, which are then used to test and compare the model’s performance. The capability dimensions include language generation, knowledge utilization, and complex reasoning, among others. Depending on the evaluation method, the assessment approaches for these capability dimensions can be classified into three categories: benchmark-based methods [81], human evaluation-based methods [82], and model evaluation-based methods. Table IV lists typical works associated with each evaluation method.
TABLE IV: Typical Evaluation Work.
Evaluation Work Capability Data Source
MMLU [81] General Human Exams
BIG-Bench [83] General Human Annotated
C-Eval [84] General Human Exams
Chatbot Arena [82] Human Interaction Human Annotated
AlpacaEval [85] Instruction Following Synthetic
MT-Bench [82] Human Interaction Human Annotated
III IIoT: Data Reservoir for LMs
III-A Sensing
Sensing technologies are fundamental to the realization of intelligent IIoT systems, enabling task recognition, environmental perception, and system optimization through diverse data modalities [86]. The performance metrics, evaluation indicators and architecture of data collection, data routing, virtualization, and other aspects have gained widespread attention [87].
  • Motion Data Inertial Measurement Unit (IMU) sensors, including accelerometers, gyroscopes, and magnetometers, are widely used for capturing motion data. Liu et al. [88] proposed the Lasagna framework, an unsupervised learning-based Human Activity Recognition (HAR) method. This framework extracts shared fundamental features of human motion in an unsupervised manner, building multi-resolution general representations for common human activities, significantly enhancing the expressiveness and utility of motion data.
  • Wireless Signals Wireless communication technologies, such as RFID, Wi-Fi, millimeter-wave, LTE, and LoRa, provide valuable signal data for various industrial applications. Li et al. [89] designed an RFID-based sensing system by attaching RFID tags to objects in clinical environments. They recorded the received signal strength from these tags and used the collected data as input to a CNN. This method enabled activity recognition involving specific objects, demonstrating significant potential in healthcare scenarios.
  • Vision Information Vision data is captured using RGB cameras, depth cameras, and Near-Infrared (NIR) image sensors. Shim et al. [90] employed NIR image sensors to monitor human activities. Unlike RGB-based vision systems, NIR images inherently lack sufficient information to reveal personal identities, effectively addressing privacy concerns in applications requiring privacy-sensitive monitoring.
  • Acoustic Signals Acoustic sensors capture sound-based data for various industrial scenarios. Mao et al. [91] introduced the DeepRange system, which addressed the limitations of traditional signal processing methods in utilizing aquatic signals for localization tasks, particularly in low Signal-to-Noise Ratio environments. They demonstrated that Deep Neural Networks could autonomously learn features from received acoustic signals to estimate distances, outperforming traditional signal processing algorithms in certain scenarios.
  • Point Cloud Data LiDAR and other 3D sensors generate high-precision point cloud data for advanced applications. Point cloud data has garnered significant attention in end-to-end autonomous driving systems [92]. For instance, high-precision point cloud data generated by LiDAR provides detailed three-dimensional environmental information, supporting reliable path planning, object detection, and environmental perception in autonomous vehicles.
III-B Data Freshness
In the IIoT environment, combining Age of Information (AoI) and Age of Incorrect Information (AoII) metrics, intelligent scheduling strategies, multi-access protocols, and sleep scheduling mechanisms effectively ensures the freshness and accuracy of information, thereby enhancing the efficiency and reliability of monitoring and control systems [93].
  • Assessing Information Freshness: AoI serves as a critical metric for quantifying freshness in state update systems by measuring the time elapsed since the destination node last received updated information [94, 95]. Unlike traditional communication delay metrics, AoI comprehensively accounts for transmission, propagation, and queuing delays. To further evaluate both freshness and accuracy, AoII [96] extends AoI by considering whether the information at the remote server aligns with the source’s actual state, providing a more holistic perspective.
  • Intelligent Scheduling Strategies: Traditional scheduling methods, such as simple priority queues or predefined rules, often fail to address the complexity and dynamism of IIoT environments. Intelligent scheduling, particularly those leveraging DRL, dynamically optimizes scheduling by learning from environmental interactions [96]. DRL excels in handling large state and action spaces while adapting to the heterogeneity of source updates and bandwidth constraints in IIoT.
  • Multi-Access Protocols and Sleep Scheduling: To reduce power consumption, various sleep scheduling strategies, such as Multi-Vacation and Start-up Threshold strategies, are applied in IIoT systems, allowing devices to enter sleep mode when idle [95]. When combined with multi-access protocols like TDMA, FDMA, and NOMA, these strategies further enhance resource allocation and network performance. Different combinations of protocols and strategies exhibit varying impacts on the system’s AoI and AoII performance.
III-C Data Augmentation
Figure 12: Data augmentation.
Data augmentation plays a critical role in enhancing the performance and robustness of LLMs by enriching the training data with synthesized or modified examples.
For instance, LLMs are particularly adept at augmenting data interpretability by generating textual descriptions for non-textual data [97]. LLMs also excel in integrating diverse industrial data sources, thereby enhancing the robustness of data processing and decision-making capabilities. For example, Hou et al. [98] utilize GPT-4 to refine hierarchical feature definitions, significantly reducing the reliance on manual annotations and improving model stability during testing phases. As shown in Fig. 12, Chen et al. enhance original images using an image-editing large model across three dimensions—time, weather, and occlusion [63].
The following sections first outline the strategies employed by LLMs for data denoising and selection.
III-C1 Data Denoise
@@ III-C2 Data Selection
III-C3 Reprogramming
Figure 13: Model reprogramming.
III-D Data Generation
Figure 14: Different types of generative models.
III-D1 Generative Models
@@
Furthermore, we include a comparison between raw and generated data, as illustrated in Fig. 15, highlighting the ability of DDPM to produce samples closely resembling the original time series data.
@@
III-D2 Generation in IIoT
We analyze the relationship between communication networks and large models (e.g., AIGC), and Table V clearly presents the commonalities and differences across these studies. The commonalities indicate that all the studies focus on optimizing AIGC services by utilizing communication networks and/or edge computing resources, highlighting the close relationship between communication networks and large models. The differences are analyzed across three dimensions: research focus, key technologies, and application scenarios. These differentiated analyses help readers gain a clearer understanding of the unique contributions of each study.
In conclusion, the integration of communication networks and large models is showing a trend of diversification. Various research directions and technical approaches provide valuable insights for the practical application of AIGC services. As research progresses, the collaborative development of communication networks and large models is expected to open up more possibilities for future intelligent applications.
TABLE V: Comparative analysis of research studies focusing on communication networks and AIGC.
Research Main Focus Main Technologies Application Scenario
EUCI [114] Inferring diffusion models for AIGC services. Model inference Resource-constrained
AIGC-IoV [115] Bridging generative AI with vehicular networks to enhance AIGC services. Network integration, model deployment Urban areas
SemCom Networks [116] Inferring service requirements and enhancing network architecture for AIGC services. Network architecture (including protocol, topology customization, etc.) Smart cities
Physical Layer [117] Enhancing physical layer security for AI-powered communication networks. Physical layer security (encryption protocols, security algorithms, etc.) Critical infrastructure
Pedestrian Detection [118] Enhancing AI-driven security systems through data augmentation. Data augmentation techniques Public safety
Mobile AIGC [111] Optimizing AIGC services for mobile networks. Network flexibility, efficiency, security, data management Mobile networks
IV IIoT: Training Ground for LMs
IV-A Pre-training
Figure 16: The schematic diagram of the Photon [119] system.
The integration of FL with GII shows considerable promise for future advancements [123]. When combined with IIoT, FL enhances application possibilities across various domains, including data sharing, task offloading, attack detection, localization, and privacy/security enhancement [124, 125]. Specifically, FL-based approaches have proven valuable in sectors such as smart healthcare, smart transportation, UAVs, smart cities, and intelligent manufacturing, where data privacy and distributed computing are critical.
IIoT devices are often widely distributed, and network quality can vary significantly, especially in cross-island federated learning scenarios. Bandwidth limitations can severely impact communication efficiency, thereby affecting training performance. Photon [119] addresses this challenge by enabling end-to-end pretraining of LLMs in low-bandwidth, globally distributed environments. Designed for collaborative training across regions and institutions, Photon outperforms centralized models by improving communication efficiency. It employs low-bandwidth optimization algorithms, such as LocalSGD, and adapts to varying client connection conditions with a dynamic local parallel mechanism, reducing communication overhead. Additionally, Photon compresses model updates using techniques like model quantization and sparse updates, significantly reducing the data transmitted per communication round. Fig. 16 illustrates the Photon system’s schematic.
IV-A1 Bandwidth Limitation and Communication Efficiency
IIoT devices are often widely distributed, and the quality of network connections can vary significantly. This is particularly evident in cross-island federated learning scenarios, where bandwidth limitations can significantly impact communication efficiency, thereby affecting training performance. To address this challenge, Photon [119], a federated learning system, focuses on end-to-end pretraining of LLMs in low-bandwidth, globally distributed environments. Its primary goal is to enable collaborative training across regions and institutions, ultimately producing models that outperform those trained via centralized approaches. To enhance communication efficiency, Photon introduces low-bandwidth optimization algorithms such as LocalSGD, and combines them with an adaptive local parallel mechanism that dynamically adjusts distributed and low-bandwidth algorithms based on client connection conditions, effectively reducing communication overhead. Moreover, Photon reduces communication frequency and compresses model updates, further decreasing communication volume. It also employs techniques like model quantization and sparse updates to compress the amount of data transmitted per communication round, significantly improving the system’s communication efficiency. Fig. 16 illustrates a schematic diagram of the Photon system.
The common goal of these studies is to reduce communication overhead in FL, thereby improving training efficiency, especially in environments with limited bandwidth and high latency. The research primarily focuses on wireless networks and IoT environments, which often face challenges such as bandwidth constraints and device heterogeneity, making the optimization of the FL communication process essential. To mitigate communication costs, many studies employ techniques such as client selection [126, 127], bandwidth allocation [126, 127, 128], and knowledge distillation [129, 130] to reduce the amount of data transferred and enhance training efficiency.
Although these studies share similar objectives and technical directions, they exhibit differences in focus and implementation. Some studies concentrate on algorithmic optimizations, such as reducing the amount of data required for model updates through knowledge distillation or optimizing bandwidth usage through joint client selection and resource allocation algorithms. In contrast, other studies emphasize communication technologies, such as clustering-based Federated Learning or utilizing multi-carrier NOMA techniques to improve bandwidth utilization and communication efficiency [131]. Additionally, the methodological approaches differ, with some studies employing optimization theory and linear programming to solve resource allocation issues [126], while others use deep reinforcement learning (DRL) to dynamically adjust scheduling strategies [127]. These differences indicate that various technical solutions are suited to different application scenarios, and selecting the appropriate approach requires considering the specific environment. We compare the relevant literature, and the results are shown in Table VI.
TABLE VI: Pre-training in Bandwidth and Communication Constrained IIoT.
Method Wireless/IoT Client Selection Bandwidth Allocation KD DRL NOMA Clustering
KD-DP [129] $\checkmark$ $\checkmark$
Fed-EC [131] $\checkmark$ $\checkmark$
FedSKD [130] $\checkmark$ $\checkmark$
J-CSBA [126] $\checkmark$ $\checkmark$ $\checkmark$
DS-BA [127] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MC-NOMA [128] $\checkmark$ $\checkmark$ $\checkmark$
IV-A2 Latency Issues
Network latency is one of the primary challenges in FL, as it can lead to delays in model parameter transmission and synchronization, significantly prolonging training time. To mitigate this issue, recent studies have proposed a variety of strategies, primarily focusing on asynchronous learning, adaptive optimization, and model compression.
To provide a comprehensive overview, we summarize these recent advancements in Table VII. The table uses the following abbreviations to describe the primary techniques employed in each study:
AA refers to asynchronous aggregation, AS stands for adaptive strategy, MC represents model compression, HM denotes heartbeat mechanism, HA indicates hierarchical aggregation, BC refers to bias correction, NS stands for neighbor selection, and DRL represents deep reinforcement learning.
First, asynchronous federated learning (AFL) has been extensively studied. Its core idea is to allow clients to update model parameters asynchronously, avoiding the constraints of global synchronization and reducing sensitivity to network latency (e.g., AS-AFL [132], MD-FEEL [133]). Asynchronous methods often optimize communication efficiency through local aggregation, neighbor selection, or hierarchical architectures, while employing bias correction techniques to address potential model update inconsistencies.
Second, adaptive optimization strategies dynamically adjust network parameters, such as model aggregation frequency, synchronization degree, and bandwidth allocation, to minimize communication delay and model loss (e.g., adaptive semi-asynchronous federated learning [134]). These strategies often integrate network conditions and device resources, optimizing client selection and resource scheduling to enhance overall system efficiency.
Additionally, model compression have emerged as an essential approach to addressing network latency. Techniques such as quantization and sparsification reduce the communication overhead of model updates, significantly lowering bandwidth requirements while maintaining model convergence (e.g., QuAsyncFL [135]). Some studies further combine these methods with hierarchical aggregation frameworks to optimize data flow and reduce cross-layer communication costs [133].
TABLE VII: Comparison of Approaches to Solving Network Latency.
Method AA AS MC HM HA BC NS DRL
AS-AFL [132] $\checkmark$ $\checkmark$ $\checkmark$
ASAFL [134] $\checkmark$ $\checkmark$
AsyDFL [136] $\checkmark$ $\checkmark$ $\checkmark$
MD-FEEL [133] $\checkmark$ $\checkmark$
QuAsyncFL [135] $\checkmark$ $\checkmark$
J-DLAL [137] $\checkmark$ $\checkmark$ $\checkmark$
Semi-AFL [138] $\checkmark$ $\checkmark$ $\checkmark$
IV-A3 Scalability
As the number of IIoT devices continues to grow, FL systems must be highly scalable to efficiently address the challenges posed by large-scale device participation in training. Recent studies have proposed various solutions, focusing primarily on hierarchical architectures, decentralized designs, clustering methods, and reputation mechanisms.
First, Hierarchical Federated Learning (HFL) has emerged as a key approach to enhancing scalability [139, 140, 141, 142]. By introducing intermediate aggregation nodes at different levels (e.g., cloud, edge, and device layers), HFL effectively reduces communication overhead and latency while improving model training efficiency. For instance, HFL methods are widely applied in edge computing and vehicular networks [142], where local aggregation at edge servers or intermediate nodes minimizes direct communication with cloud servers. These methods also incorporate clustering based on social context or device characteristics to enhance model personalization and accuracy.
Second, Decentralized Federated Learning (DFL) offers an architecture that eliminates reliance on a central server, significantly improving system robustness and fault tolerance [143, 140, 144]. By leveraging distributed topologies, DFL facilitates efficient model transmission and asynchronous updates among nodes, mitigating the risks associated with central server failures. Furthermore, optimization algorithms have been proposed to enhance model consistency in DFL, improving generalization performance in heterogeneous data and complex network environments.
Additionally, clustering-based participant selection methods have shown potential in improving FL efficiency. By grouping devices based on resource availability, data distribution, or social context information, clustering optimizes device allocation and model training processes. In Internet of Medical Things (IoMT) applications, clustering is used to select the most suitable participants, thereby enhancing system performance and reducing resource consumption [141].
Lastly, reputation mechanisms are widely employed to enhance system robustness [142]. By dynamically evaluating participant reliability and designing robust aggregation algorithms, these mechanisms effectively defend against malicious devices or intermediate nodes. For example, hierarchical architectures incorporating reputation updates dynamically adjust model aggregation weights, further improving model security and training accuracy.
Summary: To address scalability challenges posed by large-scale device participation, current research focuses on the following strategies:
  • HFL: Reducing communication overhead and improving training efficiency through intermediate aggregation nodes.
  • DFL: Eliminating reliance on central servers to enhance system robustness.
  • Clustering: Optimizing participant allocation to improve model personalization and efficiency.
  • Reputation: Dynamically evaluating participant reliability to enhance system security.
We present a comparative summary of the related studies in Table VIII.
TABLE VIII: Comparison of approaches to solving scalability.
Research Method Hierarchical FL Decentralized FL Clustering Reputation
DAHFL [145] $\checkmark$ $\checkmark$
FedLEO [143] $\checkmark$
FedUC [140] $\checkmark$ $\checkmark$ $\checkmark$
SC-HFL [141] $\checkmark$ $\checkmark$
DFedSAM [144] $\checkmark$
RoHFL [142] $\checkmark$ $\checkmark$
IV-A4 Security Concerns
Security risks in IIoT environments are relatively high, as model parameters may be susceptible to attacks during transmission, leading to data leakage or model tampering. To address this, techniques like encryption and differential privacy are employed to protect model parameters [146, 147]. Additionally, blockchain technology is proposed to enhance data security and trustworthiness, ensuring transparent, tamper-proof transmission of model parameters and thereby improving the overall security of federated learning systems. Ddeep learning methods can be applied to develop enhanced security solutions, further improving the ability to defend against potential attacks and boosting system robustness [148].
IV-B LMs as the Tuning Predictor